OpenStudy (anonymous):

1. Solve the equation for y. 4x 5y = 9 please right in fraction form

OpenStudy (rulnick):

Is it truly 4x times 5y? If so, 20xy=9 y=9/(20x)

OpenStudy (campbell_st):

the there an operator between 4x and 5y...?

OpenStudy (anonymous):

-4x-5y=9 srry

OpenStudy (rulnick):

-5y = 9 + 4x y = (9 + 4x) / (-5) y = -(9+4x)/5

OpenStudy (rulnick):

4x + 9 y = - ---------- 5

OpenStudy (anonymous):



OpenStudy (anonymous):

this is what it need to look like :)

OpenStudy (rulnick):

Yes, that's it.